# Quantifying Uncertainty in Error Consistency: Towards Reliable Behavioral Comparison of Classifiers

### v1

@@ Abstract
Benchmarking models is a key factor for the rapid progress in machine learning (ML) research. Thus, further progress depends on improving benchmarking metrics. A standard metric to measure the behavioral alignment between ML models and human observers is error consistency (EC). EC allows for more fine-grained comparisons of behavior than other metrics such as e.g. accuracy, and has been used in the influential Brain-Score benchmark to rank different DNNs by their behavioral consistency with humans. Previously, EC values have been reported without confidence intervals. However, empirically measured EC values are typically noisy—thus, without confidence intervals, valid benchmarking conclusions are problematic. Here we improve on standard EC in two ways: First, we show how to obtain confidence intervals for EC using a bootstrapping technique, allowing us to derive significance tests for EC. Second, we propose a new computational model relating the EC between two classifiers to the implicit probability that one of them copies responses from the other. This view of EC allows us to give practical guidance to scientists regarding the number of trials required for sufficiently powerful, conclusive experiments. Finally, we use our methodology to revisit popular NeuroAI-results. We find that while the general trend of behavioral differences between humans and machines holds up to scrutiny, many reported differences between deep vision models are statistically insignificant. Our methodology enables researchers to design adequately powered experiments that can reliably detect behavioral differences between models, providing a foundation for more rigorous benchmarking of behavioral alignment.
@@ 1 Introduction
Consider the following problem: Given two classifiers operating on the same domain, how should we quantify their similarity? This abstract problem has become highly practically relevant in the context of cognitive science and beyond. In cognitive science, deep neural networks (DNNs) have been proposed as models of the human visual system (Doerig et al., 2023; Kriegeskorte, 2015; Cichy and Kaiser, 2019; Kietzmann et al., 2017; Yamins et al., 2014), and thus the question arises how similarly they behave to human observers. In other domains, e.g. law or medicine, it is also crucial to know whether DNNs interpret and judge information similarly to humans. Thus, a metric is needed that can reliably quantify the degree of behavioral similarity between DNNs and human decision makers.
The standard method proposed for this purpose is error consistency (EC) (Geirhos et al., 2020), which has seen wide application in the context of human-machine comparisons, e.g. in Geirhos et al. (2021); Ollikka et al. (2024); Li et al. (2025); Parthasarathy et al. (2023). Similarity metrics based on EC have also seen application in the behavioral similarity section of the vision-branch of Brain-Score (Schrimpf et al., 2018). More recently, a variation of EC—sharing many of the relevant properties, so that our considerations apply—has been proposed as a metric to compare Large Language Models (LLMs) (Goel et al., 2025).
EC builds on Cohen’s $\kappa$ (Cohen, 1960), which considers the trial-level agreement observed for a pair of classifiers to compute a scalar similarity score, which is $1$ if the two models are behaviorally indistinguishable and $-1$ if they are as different as they could possibly be. A value of 00 indicates independence. In practice, EC is reported as a point estimate of a unit-less scalar value, based on classification decisions obtained on a single set of test stimuli, like in Figure 2a. However, since EC relies on trial-by-trial similarity, the measure is inherently noisy, as individual trials can have a large impact on the final score, especially when the total number of trials is small—as is often the case when conducting experiments with humans.
@@
  • We explain how confidence intervals for empirical EC values can be estimated on any dataset on which EC may be calculated using bootstrapping (Efron and Tibshirani, 1994).
@@
  • We apply our methods to two influential benchmarks (Model-VS-Human (Geirhos et al., 2021) and Brain-Score (Schrimpf et al., 2018)), showing that the currently available amount of human reference data is not sufficient to draw conclusions about model differences in many cases, as illustrated in Figure 1.
@@
We will begin by providing a brief introduction to EC in Section 2, to then explain how confidence intervals can be calculated in Section 3. A detailed Related Work section is provided in Appendix A since we discuss related work on measuring behavioral alignment, Error Consistency, and benchmarking.
@@ 2 Error Consistency
First, we provide a brief introduction to error consistency and introduce our notation; see Geirhos et al. (2020) for an in-depth description. EC is measured by applying Cohen’s $\kappa$ to evaluate whether two classifiers’ responses are jointly correct or incorrect—that is, whether they are consistent about when they make errors. The experimental setting in which EC can be calculated is the following: Two classifiers categorize each of $N$ samples (e.g., natural images) into one of $K$ classes (e.g., the 1,000 ImageNet classes). Each sample, $x_{i}$, has a ground-truth label $y_{i}\in\{1,...,K\}$ and and receives classifications, $\hat{y}_{i}^{(j)}\in\{1,...,K\}$, with $j\in\{1,2\}$. Whether these classifications are correct or not is coded as $r_{i}^{(j)}\in\{0,1\}$ where $r_{i}^{(j)}=1$ if and only if the $j$-th classifier responded with the true label, $\hat{y}_{i}^{(j)}=y_{i}$. This renders EC agnostic to the number of classes of the classification task because it only considers agreement on which samples are jointly classified correctly and incorrectly.
@@
However, the normalization does not render EC fully orthogonal to the accuracies of the classifiers, because Cohen’s $\kappa$ depends on the marginal distributions of classifier responses (Falotico and Quatto, 2014; Grant et al., 2016). This is easy to see in the limiting case where two classifiers have perfect accuracies ($p_{1}=p_{2}=1$): Division by zero renders EC undefined because, intuitively, error consistency can not be evaluated in the absence of errors. Moreover, if exactly one classifier has perfect accuracy, EC will be 0 irrespective of the accuracy of the other classifier. This phenomenon, which we prove in Appendix C, may be counterintuitive and can lead to estimation instabilities, see below.
@@
EC is bounded by the mismatch of the two classifiers’ accuracies (Geirhos et al., 2020). We provide an instructive visualization of these asymmetric bounds in Figure 2a. Only when the accuracies of both classifiers are equal can all values in $[-1,1]$ be achieved. This implies that EC can never be high between two classifiers if their accuracies are substantially different. In practice, one can therefore not take EC at face value: EC has to be interpreted with the accuracies of the two classifiers in mind, and is, unfortunately, strongly susceptible to both floor- and ceiling-effects. The precondition for a straightforward interpretation of EC is thus a good match of classifiers’ accuracies.
@@
Figure 2: Theoretical issues with EC. (a) The bounds on error consistency (EC) depend on the mismatch between the accuracies of the two compared classifiers. (b) Similar to (a) but with accuracies on the x and y axes. Here, each orange dot corresponds to one model-vs-human comparison and each black dot to one human-vs-human comparison in the analysis of Geirhos et al. (2021). Red dots correspond to conditions that were sampled but excluded for the analysis by the original authors. Human-vs-human comparisons have lower accuracy mismatches, so EC values are expected to be higher a priori.
@@
However, previous work has evaluated EC in conditions in which accuracies of the two classifiers were vastly different, see Figure 2b. Even without considering their joint probabilities, EC between classifiers must be low in these cases because their accuracy mismatch imposes an upper bound on the observable EC values.
@@
To formalize this dependence of EC on classifiers’ accuracy match, we present two propositions. First, when the accuracies of the two classifiers are equal, EC can best be understood as the proportion of copied responses: In a model in which the second classifier copies the responses of the first with probability $p_{copy}$ and responds independently at random otherwise, EC is related to that copy probability. In Proposition 1, we present this result as a general property of Cohen’s $\kappa$ (for relating $\hat{y}^{(1)}$ and $\hat{y}^{(2)}$) which immediately translates to the application as EC (relating $r^{(1)}$ and $r^{(2)}$), where equal marginal distributions correspond to equal accuracies.
@@
Let a classifier give responses according to some marginal distribution $\hat{y}^{(1)}\sim\tilde{P}^{(1)}$. If another classifier copies that response $\hat{y}^{(2)}=\hat{y}^{(1)}$ with probability $p_{copy}$ and otherwise draws from the same marginal distribution $\hat{y}^{(2)}\sim\tilde{P}^{(2)}$ with $\tilde{P}^{(1)}=\tilde{P}^{(2)}$, then
[…]
@@
For the joint distribution, we have $P^{(1,2)}(y,y)=P^{(1)}(y)(p_{copy}+(1-p_{copy})P^{(2)}(y))$ because the second classifier makes the same prediction as the first not only in all copy cases ($p_{copy}$), but also in those cases in which it does not copy but coincidentally makes the same prediction ($(1-p_{copy})P^{(2)}(y)$). With that, $\kappa$ simplifies to $p_{copy}$.
@@
$=\frac{\sum_{y=1}^{K}\Big{(}P^{(1)}(y)(p_{copy}+(1-p_{copy})\tilde{P}_{2}(y))-P^{(1)}(y)P^{(2)}(y)\Big{)}}{1-\sum_{y=1}^{K}P^{(1)}(y)P^{(2)}(y)}$
@@
    the (mis-)match between their accuracies (factor $f$ in Proposition 2).
Both are intuitively valid determinants of the similarity between two classifiers. We formalize the interplay between these two aspects in Proposition 2. In contrast, Safak (2020) tackled the same problem but proposed a linear scaling whereas our result reflects a more interpretable multiplicative scaling. Again, this result holds for Cohen’s $\kappa$ in general and, therefore, also when it is applied to measure EC.
@@
Let a classifier give responses according to some marginal distribution $\hat{y}^{(1)}\sim\tilde{P}^{(1)}$. If another classifier copies that response $\hat{y}^{(2)}=\hat{y}^{(1)}$ with probability $p_{copy}$ and otherwise draws from its (potentially different) marginal distribution $\hat{y}^{(2)}\sim\tilde{P}^{(2)}$, then
[…]
@@
See Appendix B for the proof. Note that the relation between EC and $p_{copy}$ is not symmetric: Whether classifier A is considered to copy from classifier B or vice versa results in different mismatch factors $f$. But note also that this factor vanishes ($f=1$) if the reference classifier, from which responses are assumed to be copied, has a uniform marginal distribution.
@@
Beyond this contribution to interpreting EC, the generative model in Proposition 2 contributes in another way: It allows simulating data to predict the stability of real data. We will heavily use the latter to give guidance to practitioners planning how many samples they should present to real human participants in their experiments.
@@ 3 Calculating confidence intervals for EC values
Whenever an aggregate measure of empirical data is reported, it is best practice to also quantify the degree of uncertainty about the final value. Historically, this has not always been done for EC, e.g. none of Geirhos et al. (2018); Ollikka et al. (2024); Li et al. (2025) report confidence intervals. We begin by outlining how confidence intervals can be obtained for existing empirical data, using a straight-forward bootstrapping approach. After that, we make use of the generative model implied by Proposition 2 to derive confidence intervals ahead of time to plan sufficiently powerful experiments.
@@ 3.1 Using existing data for bootstrapping
Assuming that we already have access to two sequences of binary responses, $r^{(1)}$ and $r^{(2)}$ of length $N$, we can obtain a measure of uncertainty simply by bootstrapping (Efron and Tibshirani, 1994), that is, by sampling $N$ elements with replacement from the sequences, $((r^{(1)}_{i},r^{(2)}_{i})\mid i\in\{1,...,N\})$, and re-calculating EC based on the sampled trials. This is established in other fields using consistency measures (McKenzie et al., 1996; Vanbelle and Albert, 2008). Doing this repeatedly yields an empirical distribution of EC values, for which we can report a 95% confidence interval by taking the central 95% of the posterior mass, i.e. the interval $[q_{0.025},q_{0.975}]$, where $q_{p}$ denotes the $p$-th quantile of the posterior distribution. Alternatively, one could report the Highest Posterior Density Interval (HDPI), which is the narrowest interval containing a certain amount (e.g. 95%) of the posterior mass.
We apply this method of estimating CIs via bootstrapping to the data from Geirhos et al. (2021) in Section 4. In their work, several human observers (typically four to six) and 52 models were evaluated on corrupted natural images, which had to be classified into sixteen classes. There are 17 different corruptions, most of which are parameterized by a scalar intensity level, leading to multiple conditions within each corruption. Each condition contains 160 to 640 images. For every model, its EC to human observers is calculated in a hierarchical fashion: First, within each condition, the ECs to all human observers are calculated and then averaged. Then, the conditions of each corruption are averaged again, before calculating the final "human-machine error consistency" as an average over corruptions. To accurately reflect this procedure, we bootstrap through the entire calculation, obtaining a posterior over the final average itself and plot 95% percentile intervals of these averages in Figure 4.
3.2 Using Proposition 2 for simulations

### v2

@@ Abstract
Benchmarking models is a key factor for the rapid progress in machine learning (ML) research. Thus, further progress depends on improving benchmarking metrics. A standard metric to measure the behavioral alignment between ML models and human observers is error consistency (EC). EC allows for more fine-grained comparisons of behavior than other metrics such as accuracy, and has been used in the influential Brain-Score benchmark to rank different DNNs by their behavioral consistency with humans. Previously, EC values have been reported without confidence intervals. However, empirically measured EC values are typically noisy—thus, without confidence intervals, valid benchmarking conclusions are problematic. Here we improve on standard EC in two ways: First, we show how to obtain confidence intervals for EC using a bootstrapping technique, allowing us to derive significance tests for EC. Second, we propose a new computational model relating the EC between two classifiers to the implicit probability that one of them copies responses from the other. This view of EC allows us to give practical guidance to scientists regarding the number of trials required for sufficiently powerful, conclusive experiments. Finally, we use our methodology to revisit popular NeuroAI-results. We find that while the general trend of behavioral differences between humans and machines holds up to scrutiny, many reported differences between deep vision models are statistically insignificant. Our methodology enables researchers to design adequately powered experiments that can reliably detect behavioral differences between models, providing a foundation for more rigorous benchmarking of behavioral alignment.
@@ 1 Introduction
Consider the following problem: Given two classifiers operating on the same domain, how should we quantify their similarity? This abstract problem has become highly practically relevant in the context of cognitive science and beyond. In cognitive science, deep neural networks (DNNs) have been proposed as models of the human visual system (doerig2023neuroconnectionist; kriegeskorte2015deep; cichy2019deep; kietzmann2017deep; yamins2014performance), and thus the question arises how similarly they behave to human observers. In other domains, e.g. law or medicine, it is also crucial to know whether DNNs interpret and judge information similarly to humans. Thus, a metric is needed that can reliably quantify the degree of behavioral similarity between DNNs and human decision makers.
The standard method proposed for this purpose is error consistency (EC) (geirhos2020beyond), which has seen wide application in the context of human-machine comparisons, e.g. in geirhos2021partial; ollikka2024humans; li2025laion; parthasarathy2023self. Similarity metrics based on EC have also seen application in the behavioral similarity section of the vision-branch of Brain-Score (schrimpf2018brainscore). More recently, a variation of EC—sharing many of the relevant properties, so that our considerations apply—has been proposed as a metric to compare Large Language Models (LLMs) (goel2025great).
EC builds on Cohen’s $\kappa$ (cohen1960coefficient), which considers the trial-level agreement observed for a pair of classifiers to compute a scalar similarity score, which is $1$ if the two models are behaviorally indistinguishable and 1-1 if they are as different as they could possibly be. A value of 0 indicates independence. In practice, EC is reported as a point estimate of a unit-less scalar value, based on classification decisions obtained on a single set of test stimuli, like in Figure˜2a. However, since EC relies on trial-by-trial similarity, the measure is inherently noisy, as individual trials can have a large impact on the final score, especially when the total number of trials is small—as is often the case when conducting experiments with humans.
@@
  • We explain how confidence intervals for empirical EC values can be estimated on any dataset on which EC may be calculated using bootstrapping (efron1994introduction).
@@
  • We apply our methods to two influential benchmarks (Model-vs-Human (geirhos2021partial) and Brain-Score (schrimpf2018brainscore)), showing that the currently available amount of human reference data is not sufficient to draw conclusions about model differences in many cases, as illustrated in Figure˜1.
@@
We will begin by providing a brief introduction to EC in Section˜2, to then explain how confidence intervals can be calculated in Section˜3. A detailed Related Work section is provided in Appendix˜A since we discuss related work on measuring behavioral alignment, Error Consistency, and benchmarking.
@@ 2 Error Consistency
First, we provide a brief introduction to error consistency and introduce our notation; see geirhos2020beyond for an in-depth description. EC is measured by applying Cohen’s $\kappa$ to evaluate whether two classifiers’ responses are jointly correct or incorrect—that is, whether they are consistent about when they make errors. The experimental setting in which EC can be calculated is the following: Two classifiers categorize each of $N$ samples (e.g., natural images) into one of $K$ classes (e.g., the 1,000 ImageNet classes). Each sample, $x_{i}$, has a ground-truth label $y_{i}\in\{1,...,K\}$ and receives classifications, $\hat{y}_{i}^{(j)}\in\{1,...,K\}$, with $j\in\{1,2\}$. Whether these classifications are correct or not is coded as $r_{i}^{(j)}\in\{0,1\}$ where $r_{i}^{(j)}=1$ if and only if the $j$-th classifier responded with the true label, $\hat{y}_{i}^{(j)}=y_{i}$. This renders EC agnostic to the number of classes of the classification task because it only considers agreement on which samples are jointly classified correctly and incorrectly.
@@
However, the normalization does not render EC fully orthogonal to the accuracies of the classifiers, because Cohen’s $\kappa$ depends on the marginal distributions of classifier responses (Falotico_2014; Grant2016). This is easy to see in the limiting case where two classifiers have perfect accuracies ($p_{1}=p_{2}=1$): Division by zero renders EC undefined because, intuitively, error consistency can not be evaluated in the absence of errors. Moreover, if exactly one classifier has perfect accuracy, EC will be 0 irrespective of the accuracy of the other classifier. This phenomenon, which we prove in Appendix˜C, may be counterintuitive and can lead to estimation instabilities, see below.
@@
EC is bounded by the mismatch of the two classifiers’ accuracies (geirhos2020beyond). We provide an instructive visualization of these asymmetric bounds in Figure˜2a. Only when the accuracies of both classifiers are equal can all values in [1,1][-1,1] be achieved. This implies that EC can never be high between two classifiers if their accuracies are substantially different. In practice, one can therefore not take EC at face value: EC has to be interpreted with the accuracies of the two classifiers in mind, and is, unfortunately, strongly susceptible to both floor- and ceiling-effects. The precondition for a straightforward interpretation of EC is thus a good match of classifiers’ accuracies.
@@
Figure 2: Theoretical issues with EC. (a) The bounds on error consistency (EC) depend on the mismatch between the accuracies of the two compared classifiers. (b) Similar to (a) but with accuracies on the x and y axes. Here, each orange dot corresponds to one model-vs-human comparison and each black dot to one human-vs-human comparison in the analysis of geirhos2021partial. Red dots correspond to conditions that were sampled but excluded for the analysis by the original authors. Human-vs-human comparisons have lower accuracy mismatches, so EC values are expected to be higher a priori.
@@
However, previous work has evaluated EC in conditions in which accuracies of the two classifiers were vastly different, see Figure˜2b. Even without considering their joint probabilities, EC between classifiers must be low in these cases because their accuracy mismatch imposes an upper bound on the observable EC values.
@@
To formalize this dependence of EC on classifiers’ accuracy match, we present two propositions. First, when the accuracies of the two classifiers are equal, EC can best be understood as the proportion of copied responses: In a model in which the second classifier copies the responses of the first with probability $p_{copy}$ and responds independently at random otherwise, EC is related to that copy probability. In Proposition˜1, we present this result as a general property of Cohen’s $\kappa$ (for relating $\hat{y}^{(1)}$ and $\hat{y}^{(2)}$) which immediately translates to the application as EC (relating $r^{(1)}$ and $r^{(2)}$), where equal marginal distributions correspond to equal accuracies.
@@
Let a classifier give responses according to some marginal distribution $\hat{y}^{(1)}\sim\tilde{P}^{(1)}$. If another classifier copies that response $\hat{y}^{(2)}=\hat{y}^{(1)}$ with probability $p_{copy}$ and otherwise draws independently from a marginal distribution $\hat{y}^{(2)}\sim\tilde{P}^{(2)}$ that is the same as the first $\tilde{P}^{(1)}=\tilde{P}^{(2)}$, then
$\text{Cohen's }\kappa=p_{copy}.$
@@
For the joint distribution, we have $P^{(1,2)}(y,y)=\tilde{P}^{(1)}(y)(p_{copy}+(1-p_{copy})\tilde{P}^{(2)}(y))$ because the second classifier makes the same prediction as the first not only in all copy cases ($p_{copy}$), but also in those cases in which it does not copy but coincidentally makes the same prediction ($(1-p_{copy})P^{(2)}(y)$). With that, $\kappa$ simplifies to $p_{copy}$.
@@
$=\frac{\sum_{y=1}^{K}\Big(P^{(1)}(y)(p_{copy}+(1-p_{copy})\tilde{P}^{2}(y))-P^{(1)}(y)P^{(2)}(y)\Big)}{1-\sum_{y=1}^{K}P^{(1)}(y)P^{(2)}(y)}$
@@
    the (mis-)match between their accuracies (factor $f$ in Proposition˜2).
Both are intuitively valid determinants of the similarity between two classifiers. We formalize the interplay between these two aspects in Proposition˜2. In contrast, safak2020min tackled the same problem but proposed a linear scaling whereas our result reflects a more interpretable multiplicative scaling. Again, this result holds for Cohen’s $\kappa$ in general and, therefore, also when it is applied to measure EC (for which $y$ would be replaced by $r$ and different marginals correspond to different accuracies).
@@
Let a classifier give correct responses according to some marginal distribution $\hat{y}^{(1)}\sim\tilde{P}^{(1)}$. If another classifier copies that response $\hat{y}^{(2)}=\hat{y}^{(1)}$ with probability $p_{copy}$ and otherwise draws independently from its (potentially different) marginal distribution $\hat{y}^{(2)}\sim\tilde{P}^{(2)}$, then
$\text{Cohen's }\kappa=p_{copy}\cdot\underbrace{\frac{1-\sum_{y}P^{(1)}(y)^{2}}{1-\sum_{y}P^{(1)}(y)P^{(2)}(y)}}_{\text{factor $f$ due to mismatching marginals}}$
@@
See Appendix˜B for the proof. Crucially, this view of Cohen’s $\kappa$ holds with generality as long as $K=2$, without limiting assumptions about the marginal distributions. Note that the relation between EC and $p_{copy}$ is not symmetric: Whether classifier A is considered to copy from classifier B or vice versa results in different mismatch factors $f$. But note also that this factor vanishes ($f=1$) if the reference classifier, from which responses are assumed to be copied, has a uniform marginal distribution.
@@
Beyond this contribution to interpreting EC, the generative model in Proposition˜2 contributes in another way: It allows simulating data to predict the stability of real data. We will heavily use the latter to give guidance to practitioners planning how many samples they should present to real human participants in their experiments.
@@ 3 Calculating confidence intervals for EC values
Whenever an aggregate measure of empirical data is reported, it is best practice to also quantify the degree of uncertainty about the final value. Historically, this has not always been done for EC, e.g. none of geirhos2018generalisation; ollikka2024humans; li2025laion report confidence intervals. While (geirhos2020beyond) provide a basic variant of confidence intervals, these were computed in a suboptimal way: They were based on the expected agreement $p_{exp}$ rather than the actual accuracies of the two classifiers, $p_{1}$ and $p_{2}$. This creates an ambiguity because, for example, ($p_{1}=0.5,p_{2}=0.5$) as well as ($p_{2}=0.5,p_{1}=0.99$) both produce the same expected agreement of $p_{exp}=0.5$ but result in different confidence interval widths. Moreover, these confidence intervals were necessarily centered around 0, because they were constructed only for independent observers. Our approach enables calculating confidence intervals for two dependent observers by making use of the observed accuracies. We begin by outlining how confidence intervals can be obtained for existing empirical data, using a straight-forward bootstrapping approach. After that, we make use of the generative model implied by Proposition˜2 to derive confidence intervals ahead of time to plan sufficiently powerful experiments.
@@ 3.1 Using existing data for bootstrapping
Assuming that we already have access to two sequences of binary responses, $r^{(1)}$ and $r^{(2)}$ of length $N$, we can obtain a measure of uncertainty simply by bootstrapping (efron1994introduction). To do so, we sample $N$ elements with uniform probability and replacement from the sequences, $((r^{(1)}_{i},r^{(2)}_{i})\mid i\in\{1,...,N\})$, and re-calculate EC based on the sampled trials. This procedure is established in other fields using consistency measures (McKenzie_1996; Vanbelle_2008). Doing this repeatedly yields an empirical distribution of EC values, for which we can report a 95% confidence interval by taking the central 95% of the posterior mass, i.e. the interval $q_{0.025},q_{0.975}]$, where $q_{p}$ denotes the $p$-th quantile of the posterior distribution. Alternatively, one could report the Highest Posterior Density Interval (HDPI), which is the narrowest interval containing a certain amount (e.g. 95%) of the posterior mass. The only parameter of this procedure is the number of bootstraps, $M$, which we typically set to $M=10,000$, see Figure˜8 for a justification of this choice. For readers unfamiliar with the bootstrap approach, we provide a more detailed description in Appendix˜F and an explicit algorithm in Algorithm˜1.
We apply this method of estimating CIs via bootstrapping to the data from geirhos2021partial in Section˜4. In their work, several human observers (typically four to six) and 52 models were evaluated on corrupted natural images, which had to be classified into sixteen classes. There are 17 different corruptions, most of which are parameterized by a scalar intensity level, leading to multiple conditions within each corruption. Each condition contains 160 to 640 images. For every model, its EC to human observers is calculated in a hierarchical fashion: First, within each condition, the ECs to all human observers are calculated and then averaged. Then, the conditions of each corruption are averaged again, before calculating the final “human-machine error consistency” as an average over corruptions. To accurately reflect this procedure, we bootstrap through the entire calculation, obtaining a posterior over the final average itself and plot 95% percentile intervals of these averages in Figure˜4.
3.2 Using Proposition˜2 for simulations